OpenStudy (anonymous):

The function for the cost of materials to make a muffin is f(x) = Three fourths x + 3, where x is the number of muffins. The function for the selling price of those muffins is g(f(x)), where g(x) = 3x + 4. Find the selling price of 20 muffins

OpenStudy (mertsj):

Selling Price of x muffins = g(f(x))=3(3/4x+3)+4=9/4x+13 Selling Price of20 muffins = 9/4(20)+13=58